# PAC-Bayesian Reinforcement Learning Trains Generalizable Policies

### v1

@@ 1 Introduction
Deploying reinforcement learning (RL) algorithms in safety-critical applications and real-world environments requires confidence that learned policies will generalize beyond training data. Recent work has highlighted the importance of rigorous generalization guarantees for supervised learning models (Pérez-Ortiz et al., 2021). In RL, this challenge is particularly acute because algorithms learn from sequential, temporally correlated trajectories that break the standard assumption of independent and identically distributed (i.i.d.) data underlying most classical generalization theory. Since RL trajectories exhibit strong temporal dependencies, where future states depend on past actions and the evolving policy, classical analyses that rely on the i.i.d. assumption, and their associated sharp concentration bounds, cannot be directly applied to provide meaningful generalization guarantees in this setting. Facing this challenge, the PAC-Bayesian framework (McAllester, 1999; Catoni, 2007; Germain et al., 2015; Guedj, 2019; Alquier, 2024) emerges as a promising solution: its mathematical structure maintaining distributions over hypotheses rather than a single model can be extended to handle temporal dependencies through appropriate concentration inequalities.
@@
Various approaches have tried to relax the independence condition, so as to allow for dependent data. Martingale-based methods (Seldin et al., 2011; 2012) extend PAC-Bayesian analysis to sequential settings by constructing martingale sequences from the data (e.g., value function errors, Bellman residuals), then applying martingale concentration inequalities such as Azuma-Hoeffding’s (Azuma, 1967) and Freedman’s (Freedman, 1975). Although these approaches are mathematically elegant, RL problems often fail to naturally yield martingales. In contrast, all RL problems possess Markov structure by design, suggesting that concentration inequalities tailored specifically for Markov chains can provide a more natural and better suited approach to this domain.
@@
In this work, we present a novel PAC-Bayesian generalization bound for RL that explicitly accounts for Markov dependencies through the chain’s mixing time. Our key technical contribution integrates a bounded-differences condition on the negative empirical return with McDiarmid-type concentration inequality for Markov chains (Paulin, 2018), yielding a bound with explicit constants and improved scaling that avoids the vacuity of previous approaches (cf. Section 2.4). We demonstrate that this new bound is not merely a theoretical curiosity by introducing PB-SAC, an actor-critic algorithm that makes practical use of the bound as a live, optimizable performance certificate. The algorithm periodically computes the numerical value of the PAC-Bayesian bound and uses it to guide learning through posterior sampling, transforming the generalization guarantee from a passive analytical tool into an active component of a learning algorithm.
@@
Thus, the work at hand proposes the following contributions. From a theoretical perspective, we develop a PAC-Bayesian bound for RL with explicit mixing-time dependence and improved scaling with realistic trajectory lengths and discount factors relative to prior work, representing a step toward non-vacuous certificates. Algorithmically, we introduce PB-SAC, the first practical algorithm that puts into action a non-vacuous PAC-Bayesian bound as a live performance certificate in modern deep RL, including key innovations for stable optimization. Empirically, we demonstrate that our bounds remain informative across continuous-control tasks while maintaining competitive performance with state-of-the-art methods. Our approach establishes the first practical PAC-Bayesian framework for certified performance in modern RL algorithms, bridging the gap between learning theory and algorithmic practice in sequential decision-making.
@@ 2 Preliminaries
For completeness, we briefly recall the reinforcement learning and statistical learning theory concepts we rely on throughout the paper. The exposition is intentionally concise—the goal is to fix notation and state the learning theory principles that underpin our results.
@@ 2.1 Reinforcement Learning
Reinforcement Learning (RL) studies how an agent learns to make sequential decisions through interaction with an environment. Formally, the environment is modeled as a (possibly unknown) Markov Decision Process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathbb{P}(s^{\prime}\mid s,a)$ the transition kernel, $R(s,a)$ the reward function bounded in $0,R_{\max}]$, and $\gamma\in(0,1)$ the discount factor. At each time $t$, the agent observes a state $S_{t}\in\mathcal{S}$, chooses an action $A_{t}\in\mathcal{A}$ according to a policy $\pi(a|s)$, receives a reward $R_{t+1}=R(S_{t},A_{t})$, and transitions to a new state $S_{t+1}\sim\mathbb{P}(\cdot\mid S_{t},A_{t})$.
The agent’s objective is to maximise the expected discounted return
$V_{\pi}(s)\;=\;\mathbb{E}_{\pi,\mathbb{P}}\!\bigl[G_{t}\mid S_{t}=s\bigr],\qquad\text{where}\qquad G_{t}\;=\;\sum_{k=0}^{\infty}\gamma^{k}\,R_{t+k+1}.$ (1)
The function $V_{\pi}(\cdot)$ is the state–value function. The optimal value function $V^{\star}(s)=\sup_{\pi}V_{\pi}(s)$ satisfies the Bellman optimality equation
@@
RL algorithms learn either directly a policy (policy–gradient and actor–critic methods (Haarnoja et al., 2018; Konda & Tsitsiklis, 1999; Sutton et al., 1999)) or an action–value function $Q_{\pi}(s,a)$ (value–based methods such as Q‐learning and its deep variants Mnih et al. (2013)). Model–free approaches dispense with an explicit model of $\mathbb{P}$, while model–based methods leverage or learn a transition model to plan.
2.2 Probably Approximately Correct (PAC) Learning Bounds
In a supervised learning task, the instance space is taken to be the product $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is the feature space and $\mathcal{Y}$ the label space; commonly $\mathcal{Y}\subseteq\mathbb{R}$ for regression, and $\mathcal{Y}\subseteq\mathbb{N}$ for classification problems. In this setting, a learning algorithm is a function that takes in a training sample $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$ and returns a prediction function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, also referred to as a hypothesis, parametrized by $\theta\in\Theta$, where $\Theta\subset\mathbb{R}^{p}$ denotes the set of all admissible parameter vectors. An unknown data distribution $\mathcal{D}$ over $\mathcal{Z}$ is postulated, with $\mathcal{D}_{\mathcal{X}}$ denoting the marginal distribution on $\mathcal{X}$, and training samples $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$ are such that each pair $(\bm{x}_{i},y_{i})\in\mathcal{Z}$ is an independent and identically distributed (i.i.d.) random draw from $\mathcal{D}$, that is, $S\sim\mathcal{D}^{\otimes m}:=\mathcal{D}\otimes\cdots\otimes\mathcal{D}$ ($m$ copies). The “quality” of a hypothesis $f_{\theta}$ is typically assessed through a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which quantifies the discrepancy between predicted and true outputs (labels). The performance of a given hypothesis is measured by its true risk, and its empirical risk on the training sample $S$, given by
$\mathcal{L}(\theta)=\underset{{(x,y)\sim\mathcal{D}}}{\mathbb{E}}\;\bigl[\ell\bigl(f_{\theta}(\bm{x}),y\bigr)\bigr],\qquad\hat{\mathcal{L}}_{S}(\theta)=\frac{1}{m}\sum_{i=1}^{m}\ell\bigl(f_{\theta}(\bm{x}_{i}),y_{i}\bigr).$
In supervised machine learning, the goal is to learn a hypothesis $f_{\theta}$ that accurately predicts labels $y\in\mathcal{Y}$ for given inputs $\bm{x}\in\mathcal{X}$. Since training is based on a dataset $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$, a central question is: how can we ensure that the learned function $f_{\theta}$ will perform well on unseen data?
@@ 2.2 Probably Approximately Correct (PAC) Learning Bounds
$\Pr_{S\sim\mathcal{D}^{m}}\bigl\{\mathcal{L}(\theta)\leq\hat{\mathcal{L}}_{S}(\theta)+\epsilon\bigr\}\;\geq\;1-\delta.$
Choosing a small $\delta$ then translates into a high-confidence bound for $\mathcal{L}(\theta)$. Concrete PAC bounds specify how large $m$ must be (or how large the gap $\epsilon$ can be) in terms of properties of the hypothesis class, e.g. VC‑dimension, Rademacher complexity, stability, compression, etc. All of the classical PAC bounds treated $f_{\theta}$ as a deterministic output of the algorithm.
@@ 2.3 PAC-Bayesian Bounds
$\underset{\theta\sim\rho}{\mathbb{E}}[\mathcal{L}(\theta)]$ $\;\leq\;\underset{\theta\sim\rho}{\mathbb{E}}[\hat{\mathcal{L}}_{S}(\theta)]\;+\;\frac{1}{\kappa}\Bigl(\mathrm{KL}(\rho\|\mu)+\ln\tfrac{1}{\delta}+\Psi_{\ell,\mu}(\kappa,n)\Bigr)$ (3)
@@
$\Psi_{\ell,\mu}(\kappa,m)=\ln\underset{\theta\sim\mu}{\mathbb{E}}\Bigl[\exp\bigl(\kappa\,\bigl(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{S}(\theta\bigr)\bigr)\Bigr].$
[…]
@@ 2.4 Previous PAC-Bayesian Bounds in RL
Early efforts to apply PAC-Bayesian theory to RL, notably the works of Fard & Pineau (2010) and Fard et al. (2012), established the framework’s viability for model selection in batch settings. However, their bounds suffered from poor scaling with the discount factor, rendering them numerically vacuous for problems with long effective horizons typical in modern RL (Tasdighi et al., 2025). Contemporary approaches have repurposed PAC-Bayesian concepts for other algorithmic goals. Recent works have used the framework to derive training objectives for deep exploration (Tasdighi et al., 2025) or as regularizers for lifelong learning (Zhang et al., 2024). While these works demonstrate the versatility of PAC-Bayesian formalism for algorithm design, they sidestep the original goal of providing tight, computable performance certificates for modern deep RL agents. This reveals a fundamental gap: although PAC-Bayesian theory holds promise for certified RL, existing approaches either yield vacuous bounds or repurpose the framework for different objectives. To the best of our knowledge, no practical algorithm leverages non-vacuous PAC-Bayesian bounds as live performance certificates within modern RL frameworks to date.
3 A new PAC-Bayesian Generalization Bound for RL
We now present our main theoretical contribution: a new PAC-Bayesian generalization bound for reinforcement learning that explicitly accounts for temporal dependencies through the mixing time of the underlying Markov chain.
@@ 3.1 Problem Setup
As outlined earlier, our objective is to establish a high‑probability PAC‑Bayes value‑error bound for a policy operating in a Markov decision process (MDP) when the training data are dependent trajectories—possibly gathered under an off‑policy algorithm. In this section, we begin by fixing notation, then present the main results; all proofs are defered to Appendix B.
@@
We assume access to a dataset $\mathfrak{D}=\{\xi^{(1)},\dots,\xi^{(T)}\}$ of $T$ trajectories (i.e., $N=HT$ transitions in total), collected using a behavior policy $\pi_{\theta}$, parameterized by $\theta\in\Theta$. The parameters $\theta$ are drawn from a distribution $\rho\in\mathcal{P}(\Theta)$, where $\Pi=\{\pi_{\theta}:\theta\in\Theta\}$ denotes the policy class. Henceforth, we write $\xi\sim\mathcal{M}$ (resp. $\mathfrak{D}\sim\mathcal{M}^{(T)}$) to denote sampling a trajectory (resp. a set $\mathfrak{D}$ of $T$ trajectories) under the environment dynamics $\mathbb{P}$, initial state distribution $\nu$, policy $\pi_{\theta}$, and reward function $R$, in order to avoid notational overload.
We define the discounted return of a trajectory and its expected value under policy $\pi_{\theta}$ as:
@@
We now define the expected (true) loss and its empirical counterpart:
$\mathcal{L}(\theta)=-\underset{\xi\sim\mathcal{M}}{\mathbb{E}}[G(\xi)]=\underset{\mathfrak{D}\sim\mathcal{M}^{(T)}}{\mathbb{E}}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)],\quad\text{where}\quad\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)=-\frac{1}{T}\sum_{j=1}^{T}G(\xi^{(j)}).$ (5)
Following the PAC-Bayesian paradigm, we endow the parameter space $\Theta$ with a prior distribution $\mu\in\mathcal{P}(\Theta)$ selected independently of data and a posterior $\rho\in\mathcal{P}(\Theta)$ chosen after observing $\mathfrak{D}$. This formalism enables reasoning about randomized policies drawn from $\rho$ with guarantees based on their divergence from $\mu$. Crucially for our analysis, changing one transition in the data results in a quantifiable bounded effect on the empirical loss defined in equation 5:
@@
$\bigl|\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\hat{\mathcal{L}}_{\bar{\mathfrak{D}}}(\theta)\bigr|\;\leq\;\sum_{h^{\prime}=1}^{H}\sum_{j^{\prime}=1}^{T}c_{(h^{\prime},j^{\prime})}\,\mathbb{I}\Bigl[\xi^{(j^{\prime})}_{h^{\prime}}\neq\bar{\xi}^{(j^{\prime})}_{h^{\prime}}\Bigr$] (6)
Intuitively, $c_{(h,j)}$ quantifies the transition‑level influence of altering the (h,j)(h,j)‑th state–action–reward tuple on the average return. A complete derivation—including a justification of why this bound covers propagation of the perturbed transition to future steps—is given in Appendix B.2. The result yields the explicit vector $c$ used in the main Theorem 3.2.
$c_{(h,t)}=\frac{\gamma^{h-1}R_{\max}}{T},\qquad\|c\|^{2}=\frac{R_{\max}^{2}}{T(1-\gamma^{2})}\bigl(1-\gamma^{2H}\bigr).$ (7)

### v2

@@ 1 Introduction
Deploying reinforcement learning (RL) algorithms in real environments and safety-critical applications requires high confidence that the learned policies will generalize beyond the training data. Although the statistical learning literature has proved rigorous generalization guarantees for supervised learning, and despite progress in generalization guarantees for deep learning models (Pérez-Ortiz et al., 2021) in this setting, some challenges prevent extending to other settings; particularly pertaining to data assumptions.
@@
In RL, an outstanding challenge stems from the fact that algorithms learn from sequential, temporally correlated trajectories that break the standard assumption of independent and identically distributed (i.i.d.) data underlying most classical generalization theory. Since RL trajectories exhibit strong temporal dependencies, where future states depend on past actions and the evolving policy, classical analyses that rely on the i.i.d. assumption, and their associated sharp concentration bounds, cannot be directly applied to provide meaningful generalization guarantees in this setting.
Facing the aforementioned challenge, the PAC-Bayesian computations framework, which uses PAC-Bayes bounds (McAllester, 1999; Catoni, 2007; Germain et al., 2015; Alquier, 2024) as optimization objectives and for computing high confidence certificates (Pérez-Ortiz et al., 2021) that hold at distribution level, emerges as a promising solution: the analysis maintaining distributions over hypotheses rather than a single model can be extended to temporally dependent data through appropriate concentration inequalities.
@@
Various approaches have tried to relax the independent data condition, to handle data with dependence structures. Martingale-based methods (Seldin et al., 2011; 2012) extend PAC-Bayesian analysis to sequential settings by constructing martingale sequences from the data (e.g., value function errors, Bellman residuals), then applying martingale concentration inequalities such as Azuma-Hoeffding’s (Azuma, 1967) and Freedman’s (Freedman, 1975). Although these approaches are mathematically elegant, RL problems do not naturally yield martingales. Since RL data possess Markov structure by design, concentration inequalities tailored specifically for Markov chains can provide a more natural and better suited approach to this domain.
@@
In this work, we present a novel PAC-Bayesian bound for RL that explicitly accounts for Markov dependencies through the chain’s mixing time [^1] The smallest number of steps required for the chain’s distribution to become nearly indistinguishable from its stationary distribution, regardless of its initial state. . Our key technical contribution integrates a bounded-differences condition on the negative empirical return with McDiarmid-type concentration inequality for Markov chains (Paulin, 2018), yielding a bound with explicit constants and improved scaling that avoids the vacuity of previous approaches (cf. Section 2.4). We demonstrate that this new bound is not merely a theoretical curiosity by introducing PB-SAC, an actor-critic algorithm that makes practical use of the bound as a live, optimizable objective. The algorithm periodically computes the numerical value of the PAC-Bayesian bound and uses it to guide learning through posterior sampling, transforming the generalization guarantee from a passive post-training evaluation tool into an active component of a learning algorithm.
@@
Thus, the work at hand proposes the following contributions. From a theoretical perspective, we develop a PAC-Bayesian bound for RL with explicit mixing-time dependence and improved scaling with realistic trajectory lengths and discount factors relative to prior work, representing a step toward RL certificates. Algorithmically, we introduce PB-SAC, the first practical algorithm that puts into action a non-vacuous PAC-Bayesian bound as an optimizable objective in modern deep RL, including key innovations for stable optimization. Empirically, we demonstrate that bound values remain informative across continuous-control tasks while maintaining competitive performance with state-of-the-art methods.
Our approach establishes the first practical PAC-Bayesian computations framework for optimization and certification in modern RL, bridging the gap between learning theory and algorithmic practice in sequential decision-making.
@@ 2 Preliminaries
In this section, we briefly recall the definitions of reinforcement learning and statistical learning theory concepts we rely on throughout the paper. The exposition in this section is intentionally concise—the goal is to fix notation and state the learning theory principles that underpin our results.
@@ 2.1 Reinforcement Learning
Reinforcement Learning (RL) studies how an agent learns to make decisions through sequential interaction with an environment. Formally, the environment is modeled as a (possibly unknown) Markov Decision Process (MDP) denoted $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},R,\gamma)$; where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathbb{P}(s^{\prime}\mid s,a)$ the transition kernel, $R(s,a)$ the reward function bounded in $0,R_{\max}]$, and $\gamma\in(0,1)$ the discount factor. At each time $t$, the agent observes a state $S_{t}\in\mathcal{S}$, chooses an action $A_{t}\in\mathcal{A}$ according to a policy $\pi(a|s)$, receives a reward $R_{t+1}=R(S_{t},A_{t})$, and transitions to a new state $S_{t+1}\sim\mathbb{P}(\cdot\mid S_{t},A_{t})$.
The learning agent’s objective is, at least in principle, to maximize the expected discounted return
$V_{\pi}(s)=\mathbb{E}_{\pi,\mathbb{P}}\!\bigl[G_{t}\mid S_{t}=s\bigr],$ (1)
where
$G_{t}=\sum_{k=0}^{\infty}\gamma^{k}\,R_{t+k+1}.$ (2)
The function $V_{\pi}(\cdot)$ is the state–value function. It is well known that the optimal value function $V^{\star}(s)=\sup_{\pi}V_{\pi}(s)$ satisfies the Bellman optimality equation
@@
RL algorithms either learn directly a policy (policy–gradient and actor–critic methods, see e.g. Haarnoja et al. (2018); Konda & Tsitsiklis (1999); Sutton et al. (1999)), or learn an action–value function $Q_{\pi}(s,a)$ in value–based methods such as Q‐learning and its deep variants Mnih et al. (2013). Model–free approaches do not use an explicit model of $\mathbb{P}$, while model–based methods leverage or learn a transition model to plan.
2.2 Probably Approximately Correct (PAC) Bounds
In a supervised learning task, the instance space is taken to be the product $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is the feature space and $\mathcal{Y}$ the label space; commonly $\mathcal{Y}\subseteq\mathbb{R}$ for regression, and $\mathcal{Y}\subseteq\mathbb{N}$ for classification problems. In this setting, a learning algorithm is a function that takes in a training sample $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$ and returns a prediction function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, also referred to as a hypothesis, parametrized by $\theta\in\Theta$, where $\Theta\subset\mathbb{R}^{p}$ denotes the set of all admissible parameter vectors. An unknown data distribution $\mathcal{D}$ over $\mathcal{Z}$ is postulated, with $\mathcal{D}_{\mathcal{X}}$ denoting the marginal distribution on $\mathcal{X}$, and training samples $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$ are such that each pair $(\bm{x}_{i},y_{i})\in\mathcal{Z}$ is an independent and identically distributed (i.i.d.) random draw from $\mathcal{D}$, that is, $S\sim\mathcal{D}^{\otimes m}:=\mathcal{D}\otimes\cdots\otimes\mathcal{D}$ ($m$ copies). The “quality” of a hypothesis $f_{\theta}$ is typically assessed through a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which quantifies the discrepancy between predicted and true outputs (labels). Two important functionals associated to a hypothesis are its true risk on the distribution $\mathcal{D}$, and its empirical risk on the sample $S$, which are respectively given by
$\mathcal{L}(\theta)$ $=\underset{{(x,y)\sim\mathcal{D}}}{\mathbb{E}}\bigl[\ell\bigl(f_{\theta}(\bm{x}),y\bigr)\bigr],$ (4)
$\hat{\mathcal{L}}_{S}(\theta)$ $=\frac{1}{m}\sum_{i=1}^{m}\ell\bigl(f_{\theta}(\bm{x}_{i}),y_{i}\bigr).$ (5)
In supervised machine learning, the goal is to learn a hypothesis $f_{\theta}$ that accurately predicts labels $y\in\mathcal{Y}$ for given inputs $\bm{x}\in\mathcal{X}$. Since training is based on a finite dataset, a central question is: how can we ensure that the learned function $f_{\theta}$ will perform well on unseen data?
@@ 2.2 Probably Approximately Correct (PAC) Bounds
$\Pr_{S\sim\mathcal{D}^{m}}\bigl\{\mathcal{L}(\theta)\leq\hat{\mathcal{L}}_{S}(\theta)+\epsilon\bigr\}\geq 1-\delta.$
Choosing a small $\delta$ then translates into a high-confidence bound for $\mathcal{L}(\theta)$. Concrete PAC bounds specify how large $m$ must be (or how large the gap $\epsilon$ can be) in terms of properties of the hypothesis class, e.g. VC‑dimension, Rademacher complexity, stability, compression, etc.
All of the classical PAC bounds treated $f_{\theta}$ as a deterministic output of the algorithm.
@@ 2.3 PAC-Bayesian Bounds
$\underset{\theta\sim\rho}{\mathbb{E}}[\mathcal{L}(\theta)]$ $\leq\underset{\theta\sim\rho}{\mathbb{E}}[\hat{\mathcal{L}}_{S}(\theta)]+\frac{1}{\kappa}\Bigl(\mathrm{KL}(\rho\|\mu)+\ln\tfrac{1}{\delta}+\Psi_{\ell,\mu}(\kappa,m)\Bigr)$ (6)
@@
$\Psi_{\ell,\mu}(\kappa,m)=\ln\underset{\theta\sim\mu}{\mathbb{E}}\;\underset{S\sim\mathcal{D}^{\otimes m}}{\mathbb{E}}\Bigl[\exp\bigl(\kappa\,\bigl(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{S}(\theta\bigr)\bigr)\Bigr].$
Compared with classical PAC guarantees, PAC‑Bayes offers two advantages that are critical for reinforcement learning: (1) Data‑dependent priors (Parrado-Hernández et al., 2012) —when $\mu$ can itself depend on previous data, e.g. earlier tasks or behavioural trajectories (Zhang et al., 2025), the bound adapts to the knowledge already acquired, tightening $\mathrm{KL}(\rho\|\mu)$; and (2) Fine‑grained control via $\Psi$ —by tailoring the concentration inequality used to upper‑bound $\Psi$ one can incorporate dependence structures such as martingales (Seldin et al., 2012; 2011), $\beta$‑mixing (Ralaivola et al., 2009; Abélès et al., 2025) sequences or Markov chains (Fard et al., 2011; Tasdighi et al., 2025). The latter is exactly the scenario in which RL trajectories are collected.
@@ 2.4 Previous PAC-Bayesian Bounds in RL
Early efforts to apply PAC-Bayesian theory to RL, notably the works of Fard & Pineau (2010) and Fard et al. (2011), established the framework’s viability for model selection in batch settings. However, their bounds suffered from poor scaling with the discount factor, rendering them numerically vacuous for problems with long effective horizons typical in modern RL (Tasdighi et al., 2025). Contemporary approaches have repurposed PAC-Bayes bounds for other algorithmic goals. Recent works have used them to derive training objectives for deep exploration (Tasdighi et al., 2025) or as regularizers for lifelong learning (Zhang et al., 2025). While these works demonstrate the versatility of PAC-Bayesian computations for algorithm design, they sidestep the original goal of providing tight, computable certificates for modern deep RL agents. This reveals a fundamental gap: although PAC-Bayesian theory holds promise for certified RL, existing approaches either yield vacuous bounds or repurpose them for different objectives. To the best of our knowledge, no practical algorithm has leveraged non-vacuous PAC-Bayesian bounds as live performance certificates within modern RL frameworks to date.
3 A new PAC-Bayesian Bound for RL
We now present our main theoretical contribution: a new PAC-Bayes generalization bound for RL that explicitly accounts for temporal dependencies through the mixing time of the underlying Markov chain.
@@ 3.1 Problem Setup
As outlined earlier, our objective is to establish a high‑probability PAC‑Bayes value‑error bound for a policy operating in a Markov decision process (MDP) when the training data are dependent trajectories—possibly gathered under an off‑policy algorithm. In this section, we begin by fixing notation and then present the main results; all proofs are deferred to Appendix B.
@@
We assume access to a dataset $\mathfrak{D}=\{\xi^{(1)},\dots,\xi^{(T)}\}$ of $T$ trajectories (i.e., $N=HT$ transitions in total), collected using a behavior policy $\pi_{b}$. An evaluation policy $\pi_{\theta}$ is parameterized by $\theta\in\Theta$, the parameters $\theta$ are drawn from a distribution $\rho\in\mathcal{P}(\Theta)$, where $\Pi=\{\pi_{\theta}:\theta\in\Theta\}$ denotes the policy class. Henceforth, we write $\xi\sim\mathcal{M}$ (resp. $\mathfrak{D}\sim\mathcal{M}^{(T)}$) to denote sampling a trajectory (resp. a set $\mathfrak{D}$ of $T$ trajectories) under the environment dynamics $\mathbb{P}$, initial state distribution $\nu$, policy $\pi_{b}$, and reward function $R$, in order to avoid notational overload.
We define the discounted return of a trajectory $\xi$ and the value of policy $\pi_{\theta}$ respectively as:
@@
We now define the expected (true) loss and its empirical counterpart using importance sampling to account for the off-policy data distribution:
$\mathcal{L}(\theta)=-\underset{\xi\sim\mathcal{M}}{\mathbb{E}}\left[\frac{\pi_{\theta}(\xi)}{\pi_{b}(\xi)}G(\xi)\right]=-\underset{\mathfrak{D}\sim\mathcal{M}^{(T)}}{\mathbb{E}}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)],$ (8)
$\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)=-\frac{1}{T}\sum_{j=1}^{T}w_{j}(\theta)G(\xi^{(j)}),$ (9)
where $w_{j}(\theta)=\frac{\pi_{\theta}(\xi^{(j)})}{\pi_{b}(\xi^{(j)})}$ is the likelihood ratio of trajectory $\xi^{(j)}\sim\pi_{b}$. To ensure theoretical stability, we assume the weighted rewards are bounded such that the weighted return remains within the range consistent with $R_{\max}$ (e.g., via clipping large importance weights, a standard practice in off-policy evaluation).
Following PAC-Bayesian folklore, we endow the parameter space $\Theta$ with a distribution $\mu\in\mathcal{P}(\Theta)$ playing the role of a prior selected independently of data (or partial dependence on it, cf. Parrado-Hernández et al. (2012); Pérez-Ortiz et al. (2021)), and a posterior $\rho\in\mathcal{P}(\Theta)$ chosen after observing $\mathfrak{D}$. This formalism enables reasoning about randomized policies drawn from $\rho$ with guarantees based on their divergence from $\mu$. Crucially for our analysis, changing one transition in the data results in a quantifiable bounded effect on the empirical loss defined in equation 8:
@@
$\hskip-5.0pt\bigl|\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\hat{\mathcal{L}}_{\bar{\mathfrak{D}}}(\theta)\bigr|\leq\sum_{h^{\prime}=1}^{H}\sum_{j^{\prime}=1}^{T}c_{(h^{\prime},j^{\prime})}\,\mathbb{I}\Bigl[\xi^{(j^{\prime})}_{h^{\prime}}\neq\bar{\xi}^{(j^{\prime})}_{h^{\prime}}\Bigr$] (10)
Intuitively, $c_{(h,j)}$ quantifies the transition‑level influence of altering the (h,j)(h,j)‑th state–action–reward tuple on the average return. A complete derivation—including a justification of why this bound covers propagation of the perturbed transition to future steps—is given in Appendix B.2. The result yields the explicit vector $c$ used in Theorem 3.2, namely
$c_{(h,j)}=\frac{\gamma^{h-1}R_{\max}}{T},$ (11)
and its norm
$\|c\|^{2}=\frac{R_{\max}^{2}}{T(1-\gamma^{2})}\bigl(1-\gamma^{2H}\bigr).$ (12)